Can someone please walk through the steps on getting the final structure (boxed towards top)?7 IR Spectrum (liquid film) groups: -- 1712 12771254 c=0 4000 3000 1200 800 2000 1600 V (cm) ;C=0 1001 121 Mass Spectrum 80 6
Answer #1


IR: 3000 cm*: C-H(=C-H) Stretching IR: 1712 cm-1: C=0 Stretching IR: 1600 cm: C=C Stretching IR: 1277 cm?: C-o Stretching (